OpenStudy (anonymous):

Can someone help me with rationalizing the denominator? 4√6/√30

OpenStudy (jdoe0001):

what do you mean by rationalizing?

OpenStudy (jdoe0001):

hmm, ok

OpenStudy (anonymous):

to remove the radical

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

why not just multiply the fraction by 1? \( \cfrac{\sqrt{30}}{\sqrt{30}}\)

OpenStudy (anonymous):

okay. so 4√6/√30 x √30/√30 equals?? 4√180

OpenStudy (anonymous):

somebody help me so i can understand it at least?

OpenStudy (jdoe0001):

well 4√180 is correct, for the numerator, what do you get for the denominator?

OpenStudy (anonymous):

30

OpenStudy (jdoe0001):

yep

OpenStudy (jdoe0001):

so, the root is gone from the denominator you can of course, simplify it further, by say \(\sqrt{180} = \sqrt{6^2\times 5} = 4\times 6 \sqrt{5} \)

OpenStudy (jdoe0001):

and divide by the denominator, but your answer thus far is ok

OpenStudy (anonymous):

okay thanks im starting to understand it better